OpenStudy (anonymous):

At New York City's LaGuardia airport, one parking garage charges a onetime transportaton fee of $2.75 plus $22 for indoor parking per day. How many days has a car been parked at this location if the parking charge is $134.75?

OpenStudy (texaschic101):

x stands for number of days 22 + 2.75x = 134.75 2.75x = 134.75 - 22 2.75x = 112.75 2.75x/2.75 = 112.75/2.75 x = 41 check.. 22 + 2.75x = 134.75 22 + 2.75(41) = 134.75 22 + 112.75 = 134.75 134.75 = 134.75 (correct) ANSWER : The car was parked 41 days and was charged $ 134.75.